# IceCube Collaboration

### v1

@@ IV.3 Fit of inelasticity distribution to data
The analysis is performed using a forward folding method [11, 21] where the final level simulation is weighted at each combination of parameters tested, and is subsequently used to populate the histogram that is compared to data. We employ an optimization algorithm to select the values of the parameters that will be tested and minimize the function
$-\log\mathcal{L}=\sum_{i=1}^{N_{bins}}(k_{i}\log\mu_{i}-\mu_{i})-\sum_{j=0}^{M_{priors}}\frac{(p_{j}-\hat{p_{j}})^{2}}{2\sigma^{2}_{j}}$ (7)
where the sum runs over all the bins of the histogram $N_{bins}$, $\mu_{i}$ stands for the number of expected events in the $i$-th bin, which depends on the combination of parameters chosen, and $k_{i}$ is the number of observed events. The second term, where the sum is over the number of parameters with Gaussian (half-Gaussian) priors $M_{\mathrm{priors}}$, introduces prior knowledge on some of the fit parameters $p_{j}$ ($j=0,...,M_{\mathrm{priors}}$), penalizing deviations from their best estimates $\hat{p_{j}}$ in terms of their variance $\sigma^{2}_{j}$.
The neutrino simulation was produced using the GENIE Monte-Carlo generator [17] and weighted according to the expected atmospheric neutrino flux. Additional weighting factors were introduced for neutrinos and antineutrinos to produce flat $\frac{dp}{dy}$ distributions, which could then be modified using Eq. 3 for all events in the same $E_{\mathrm{reco}}$ bin. Reweighting all events to the same parametrized distribution means that we will be measuring flux-averaged $\frac{dp}{dy}$ as shown in Eq. 6.
We use the HKKMS’15 atmospheric neutrino flux model [18], and include the effects of neutrino oscillations using the two-flavor approximation ($\nu_{\mu}\rightarrow\nu_{\tau}$) and the parameters reported in Ref. [22], which is sufficient to describe the data since the low energy threshold of this study is $E_{\mathrm{reco}}$=100 GeV. Atmospheric muons were produced using MuonGun, weighted to follow the expected flux from Refs. [23, 24].
Group Parameter Best fit value Nominal value Prior
@@
Rel. eff. $p_{0}$0.5991 -0.2674 Gaussian ($\pm$ 0.6)
Rel. eff. $p_{1}$0.02251 -0.04206 Gaussian ($\pm$ 0.12)
@@
Table 2: Summary of the best fit values and settings for both the physics and nuisance parameters used in the analysis. All physics parameters are presented with $1\sigma$ uncertainties calculated using Wilks’ theorem. The $\langle y\rangle$ parameters for each energy bin include an additional uncertainty due to limited MC statistics, as described in Sec. IV.4.
@@ IV.4 Systematic uncertainties impacting the study
The impacts of all the sources of systematic uncertainties that affect this analysis were introduced as correction factors that modify the weight of simulated events. By changing the weights, these factors modify the expectation of all the bins simultaneously, producing a distinct signature in the $E_{\mathrm{reco}}$ vs $y_{\mathrm{reco}}$ space. These signatures differ from the ones produced by changes in the inelasticity distribution, and therefore allow us to perform the measurement accurately, notwithstanding the uncertainties.
@@
The impact of ice and detector properties are obtained by simulating MC sets with different input values, and parameterizing the changes they produce in the bin counts of the analysis histogram. The flux normalization is introduced as a global scaling factor, while the energy dependence is modeled as a power-law modification with a spectral index $E^{\Delta\gamma}$. The impact of the refrozen ice columns is encoded in two parameters that change the angular acceptance of module. The details of the implementation of these parameters can be found in Ref. [11].
Table 2 presents the parameters fitted in the study, including their nominal expectation and the prior knowledge, when applicable. The normalization is given an uniform prior, while the other systematic parameters that encode sources of uncertainty get a weak Gaussian prior. Additional information on the choice of priors is given in Ref. [11].
The limited Monte-Carlo statistics introduce fluctuations in the selection efficiency for neutrinos and antineutrinos, which can potentially lead to a small bias in the fitted mean inelasticity. To account for this, we include an additional error term when reporting the $\langle y\rangle$ results for a given energy bin, such that
$\langle y\rangle=\langle y\rangle_{\mathrm{best~{}fit}}\pm\sigma_{\mathrm{Wilks^{\prime}~{}th.}}\pm\Delta_{\mathrm{MC~{}stat.}}$ (8)
where $\sigma_{\mathrm{Wilks^{\prime}~{}th.}}$ is the $1\sigma$ interval obtained from the likelihood profile using the Wilks’ theorem, and $\Delta_{\mathrm{MC~{}stat.}}$ is the bias in $\langle y\rangle$ obtained using results of an inject-recover test to a sample where original GENIE inelasticity distribution was not modified by reweighting. The sample was fitted using the parametrization in Eq. 3 and the best fit $\langle y\rangle$ value for each bin was compared to the expectation for the same energy bin, calculated using generator-level simulation, with the difference being assigned to the $\Delta_{\mathrm{MC~{}stat.}}$ term.
@@ V Results
We applied the analysis to data from the 2011-2021 seasons, which correspond to 9.28 years of detector livetime. The post-fit Monte-Carlo is in good agreement with the data, with a $p$-value of 9.5%. Figure 8 shows this for $y_{\mathrm{reco}}$ in each of the three $E_{\mathrm{reco}}$ bins. The mean inelasticity is compatible with a constant value $\langle y\rangle$ = 0.45 across all energies, with a $\lambda$ parameter that differs for the lowest energies. The values for all the fit parameters are shown in Table 2. The top panel of Figure 9 shows two-dimensional confidence regions for $\langle y\rangle$ and $\log_{10}\lambda$ in each $E_{\mathrm{reco}}$ bin. The bottom panel on Figure 9 illustrates $\frac{dp}{dy}$ distributions corresponding to the best fit, as well as sampled points within each confidence region. We also show the expectation for our sample using both GENIE [17] and CSMS [27] cross section models, with the HKKMS’15 atmospheric neutrino flux model [18]. The inelasticity distributions show slight differences in shape, with our fit preferring a relatively lower contribution at low $y$, and a relatively higher contribution at high $y$.
Figure 9: Top: Likelihood space of the physical parameters $(\langle y\rangle,\log_{10}\lambda)$ with contours that denote the uncertainty at 1, 2 or 3 $\sigma$. Bottom: Inelasticity distribution shapes as function of inelasticity that results from the fit, also showing the curves that correspond to the points within each confidence region displayed on the top panel in corresponding colors. The grey lines in the background show a representative sample of the values allowed during fit that are beyond 3 $\sigma$. The figure also includes predictions that result from the combination of GENIE (H2O) or CSMS (isoscalar) cross section and the HKKMS’15 atmospheric neutrino flux calculation.
@@
Figure 12 shows the mean inelasticity compared with results obtained using higher energy interactions in IceCube [7]. Only a subset of the models tested are shown. It important to note that the mean inelasticity prediction is dependent on the event selection, so the results are not directly comparable in detail, but the general trend is consistent.
@@
Figure 12: Measured mean inelasticity $\langle y\rangle$ as a function of neutrino energy is compared to the GENIE 2.12.8 $+$ HKKMS 2015 prediction for our event sample. The IceCube 1 TeV-1 PeV inelasticity result [7] and the CSMS [27] $+$ HKKMS 2015 prediction for the corresponding sample are also shown. The highlighted energy range below 245 GeV is where measurements of differential cross section have been made by NuTeV/CCFR [32] and CDHSW [33].
The mean inelasticity can also be interpreted as a measurement of the ratio of neutrinos to antineutrinos in the atmospheric neutrino flux, if we assume a specific cross section. This comparison is shown in Fig. 13, together with the neutrino-antineutrino fraction expected from the same combinations of models discussed before. The spread of the predictions is significantly smaller than the error bars of our measurement.
@@
VI Discussion
In this study we used the observed energy in track-like and cascade-like byproducts of neutrino candidate events in IceCube to measure the inelasticity of $\nu-$nucleon interactions. We focus on a highly pure sample of $\nu_{\mu}$-CC events between 80 GeV and 560 GeV that interact inside the DeepCore volume. This is the first time that such details of the differential neutrino cross section have been measured at energies between 245 GeV and 1 TeV.
The cross section as a function of inelasticity was parameterized as before [7], and we measured the the two parameters that define its shape $\langle y\rangle$ and $\lambda$. We then compared the mean inelasticity $\langle y\rangle$ obtained with the expectation from three cross section models and two atmospheric neutrino fluxes.
@@ Acknowledgements.
We thank Yuri Onishchuk for contributions to the early phase of this work. The authors gratefully acknowledge the support from the following agencies and institutions: USA – U.S. National Science Foundation-Office of Polar Programs, U.S. National Science Foundation-Physics Division, U.S. National Science Foundation-EPSCoR, U.S. National Science Foundation-Office of Advanced Cyberinfrastructure, Wisconsin Alumni Research Foundation, Center for High Throughput Computing (CHTC) at the University of Wisconsin–Madison, Open Science Grid (OSG), Partnership to Advance Throughput Computing (PATh), Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS), Frontera and Ranch computing project at the Texas Advanced Computing Center, U.S. Department of Energy-National Energy Research Scientific Computing Center, Particle astrophysics research computing center at the University of Maryland, Institute for Cyber-Enabled Research at Michigan State University, Astroparticle physics computational facility at Marquette University, NVIDIA Corporation, and Google Cloud Platform; Belgium – Funds for Scientific Research (FRS-FNRS and FWO), FWO Odysseus and Big Science programmes, and Belgian Federal Science Policy Office (Belspo); Germany – Bundesministerium für Bildung und Forschung (BMBF), Deutsche Forschungsgemeinschaft (DFG), Helmholtz Alliance for Astroparticle Physics (HAP), Initiative and Networking Fund of the Helmholtz Association, Deutsches Elektronen Synchrotron (DESY), and High Performance Computing cluster of the RWTH Aachen; Sweden – Swedish Research Council, Swedish Polar Research Secretariat, Swedish National Infrastructure for Computing (SNIC), and Knut and Alice Wallenberg Foundation; European Union – EGI Advanced Computing for research; Australia – Australian Research Council; Canada – Natural Sciences and Engineering Research Council of Canada, Calcul Québec, Compute Ontario, Canada Foundation for Innovation, WestGrid, Digital Research Alliance of Canada, and Arthur B. McDonald Canadian Astroparticle Physics Research Institute; Denmark – Villum Fonden, Carlsberg Foundation, and European Commission; New Zealand – Marsden Fund; Japan – Japan Society for Promotion of Science (JSPS) and Institute for Global Prominent Research (IGPR) of Chiba University; Korea – National Research Foundation of Korea (NRF); Switzerland – Swiss National Science Foundation (SNSF).
@@ Appendix A Limitations of the parametrization
The physical interpretation of Eq. 3 requires $y=[0,1]$, which means that $0<$$\langle y\rangle$ $<1$, and $\frac{dp}{dy}$ must be positive, which in turn requires that $\lambda>0$ and $\epsilon>-1$. The allowed physical space as well as these limits are shown in Fig. 14.
A special case arises when $\lambda<1$, because in that case $\lim_{y\to 0}\frac{dp}{dy}=\infty$. Because of the finite number of events in our data, and also because of presence of kinematic constraints on events with very low inelasticity, the fit might not be able to detect a sharp rise in the distribution near $y=0$. Since there is no data in this region, the integral of the distribution diverges from $1$, thus violating the probability distribution function normalization requirement. This has two undesirable consequences in our analysis. First, this means that the $\langle y\rangle$ variable in Eq. (3) loses it physical interpretation as the mean inelasticity. The second outcome is that breaking of the normalization requirement can cause the minimizer to use $\frac{dp}{dy}$ parameters to modify normalization of the data, which will absorb part of the systematic uncertainties in the study, affecting the correctness of the measurement.
To avert this problem, we adjusted the simulation to data in $\log_{10}\lambda$ instead of $\lambda$ and put a weak half-Gaussian prior on negative $\log_{10}\lambda$ with $\sigma(\log_{10}\lambda)=0.4$ on the negative side. Then, if the data prefers $\lambda<1$, the fit can still allow it with minimal bias, but if this is not the case, the prior will prevent the minimizer from getting stuck in local minima, which is especially relevant when performing a scan to determine confidence regions. The correct behavior of the fit in these cases was confirmed with Monte-Carlo tests.

### v2

@@ IV.3 Fit of inelasticity distribution to data
The analysis is performed using a forward folding method [11, 21] where the final level simulation is weighted at each combination of parameters tested and is subsequently used to populate the histogram that is compared to data. We employ an optimization algorithm to select the values of the parameters that will be tested and minimize the function
$\log\mathcal{L}=\sum_{i=1}^{N_{bins}}(k_{i}\log\mu_{i}-\mu_{i})-\sum_{j=0}^{M_{priors}}\frac{(p_{j}-\hat{p_{j}})^{2}}{2\sigma^{2}_{j}}$ (7)
where the sum runs over all the bins of the histogram $N_{bins}$, $\mu_{i}$ stands for the number of expected events in the $i$th bin, which depends on the combination of parameters chosen, and $k_{i}$ is the number of observed events. The second term, where the sum is over the number of parameters with Gaussian (half-Gaussian) priors $M_{\mathrm{priors}}$, introduces prior knowledge on some of the fit parameters $p_{j}$ ($j=0,...,M_{\mathrm{priors}}$), penalizing deviations from their best estimates $\hat{p_{j}}$ in terms of their variance $\sigma^{2}_{j}$.
The neutrino simulation was produced using the genie Monte Carlo generator [17] and weighted according to the expected atmospheric neutrino flux. Additional weighting factors were introduced for neutrinos and antineutrinos to produce flat $\frac{dp}{dy}$ distributions, which could then be modified using Eq. (3) for all events in the same $E_{\mathrm{reco}}$ bin. Reweighting all events to the same parametrized distribution means that we will be extracting flux-averaged $\frac{dp}{dy}$ as shown in Eq. (6).
We use the HKKMS’15 atmospheric neutrino flux model [18] and include the effects of neutrino oscillations using the two-flavor approximation ($\nu_{\mu}\rightarrow\nu_{\tau}$) and the parameters reported in Ref. [22], which is sufficient to describe the data since the low-energy threshold of this study is $E_{\mathrm{reco}}$=100 GeV. Atmospheric muons were produced using MuonGun, weighted to follow the expected flux from Refs. [23, 24].
Group Parameter Best-fit value Nominal value Prior
@@
Relative efficiency $p_{0}$0.5991 -0.2674 Gaussian ($\pm$ 0.6)
Relative efficiency $p_{1}$0.02251 -0.04206 Gaussian ($\pm$ 0.12)
@@
Table 2: Summary of the best-fit values and settings for both the physics and nuisance parameters used in the analysis. All physics parameters are presented with $1\sigma$ uncertainties calculated using Wilks’ theorem. The $\langle y\rangle$ parameters for each energy bin include an additional uncertainty due to limited MC statistics, as described in Sec. IV.4.
@@ IV.4 Systematic uncertainties impacting the study
The impacts of all the sources of systematic uncertainties that affect this analysis were introduced as correction factors that modify the weight of simulated events. By changing the weights, these factors modify the expectation of all the bins simultaneously, producing a distinct signature in the $E_{\mathrm{reco}}$ vs $y_{\mathrm{reco}}$ space. These signatures differ from the ones produced by changes in the inelasticity distribution and therefore allow us to perform the measurement accurately, notwithstanding the uncertainties.
@@
The impact of ice and detector properties are obtained by simulating Monte Carlo sets with different input values and parametrizing the changes they produce in the bin counts of the analysis histogram. The five quantities used to parametrize said properties are scaling factors for DOM efficiency, ice absorption and scattering coefficients, and two parameters (relative efficiency $p_{0}$ and $p_{1}$), which were obtained using a principle component analysis of the DOM angular acceptance calibration data and are used to modify angular acceptance of the modules. The flux normalization is introduced as a global scaling factor, while the energy dependence is modeled as a power-law modification with a spectral index $E^{\Delta\gamma}$. The details of the implementation of these parameters can be found in Ref. [11].
Table 2 presents the parameters fitted in the study, including their nominal expectation and the prior knowledge, when applicable. The normalization is given a uniform prior, while the other systematic parameters that encode sources of uncertainty get a weak Gaussian prior. Additional information on the choice of priors is given in Ref. [11].
The limited Monte Carlo statistics introduce fluctuations in the selection efficiency for neutrinos and antineutrinos, which can potentially lead to a small bias in the fitted mean inelasticity. To account for this, we include an additional error term when reporting the $\langle y\rangle$ results for a given energy bin, such that
$\langle y\rangle=\langle y\rangle_{\mathrm{best~{}fit}}\pm\sigma_{\mathrm{Wilks^{\prime}~{}th.}}\pm\Delta_{\mathrm{MC~{}stat.}},$ (8)
where $\sigma_{\mathrm{Wilks^{\prime}~{}th.}}$ is the $1\sigma$ interval obtained from the likelihood profile using the Wilks’ theorem, and $\Delta_{\mathrm{MC~{}stat.}}$ is the bias in $\langle y\rangle$ obtained using results of an inject-recover test to a sample where original genie inelasticity distribution was not modified by reweighting. The sample was fitted using the parametrization in Eq. (3), and the best-fit $\langle y\rangle$ value for each bin was compared to the expectation for the same energy bin, calculated using generator-level simulation, with the difference being assigned to the $\Delta_{\mathrm{MC~{}stat.}}$ term. Values of the $\Delta_{\mathrm{MC~{}stat.}}$ for each of the energy bins is shown in Table 2.
@@ V Results
We applied the analysis to data from the 2011–2021 seasons, which correspond to 9.28 years of detector livetime. The post-fit Monte Carlo is in good agreement with the data, with a $p$-value of 9.5%. Figure 8 shows this for $y_{\mathrm{reco}}$ in each of the three $E_{\mathrm{reco}}$ bins. The mean inelasticity is compatible with a constant value $\langle y\rangle$ = 0.45 across all energies, with a $\lambda$ parameter that differs for the lowest energies. The values for all the fit parameters are shown in Table 2. The top panel of Fig. 9 shows two-dimensional confidence regions for $\langle y\rangle$ and $\log_{10}\lambda$ in each $E_{\mathrm{reco}}$ bin. The bottom panel on Fig. 9 illustrates $\frac{dp}{dy}$ distributions corresponding to the best fit as well as sampled points within each confidence region. We also show the expectation for our sample using both genie [17] and Cooper-Sarkar–Mertsch–Sarkar (CSMS) [27] cross section models, with the HKKMS’15 atmospheric neutrino flux model [18]. The inelasticity distributions show slight differences in shape, with our fit preferring a relatively lower contribution at low $y$ and a relatively higher contribution at medium and high $y$ values.
Figure 9: Top: likelihood space of the physical parameters $(\langle y\rangle,\log_{10}\lambda)$ with contours that denote the uncertainty at 1, 2, or 3 $\sigma$. Bottom: inelasticity distribution shapes as function of inelasticity that results from the fit, also showing the curves that correspond to the points within each confidence region displayed on the top panel in corresponding colors. The gray lines in the background show a representative sample of the values allowed during fit that are beyond 3 $\sigma$. The figure also includes predictions that result from the combination of genie or CSMS cross section (both calculated for water) and the HKKMS’15 atmospheric neutrino flux calculation.
@@
Figure 12 shows the mean inelasticity compared with results obtained using higher-energy interactions in IceCube [7]. Only a subset of the models tested are shown. It is important to note that the mean inelasticity prediction is dependent on the event selection, so the results are not directly comparable in detail, but the general trend is consistent.
@@
Figure 12: Measured mean inelasticity $\langle y\rangle$ as a function of neutrino energy is compared to the genie 2.12.8 $+$ HKKMS 2015 prediction for our event sample. The IceCube 1 TeV–1 PeV inelasticity result [7] and the CSMS [27] $+$ HKKMS 2015 prediction for the corresponding sample are also shown. The highlighted energy range below 245 GeV is where measurements of differential cross section have been made by NuTeV/CCFR [32] and CDHSW [33].
The mean inelasticity can also be interpreted as a measure of the ratio of neutrinos to antineutrinos in the atmospheric neutrino flux, if we assume a specific cross section. This comparison is shown in Fig. 13, together with the neutrino-antineutrino fraction expected from the same combinations of models discussed before. The spread of the predictions is significantly smaller than the error bars of our result.
@@
VI Conclusions
In this study, we used the observed energy in tracklike and cascadelike byproducts of neutrino candidate events in IceCube to extract the inelasticity of $\nu$-nucleon interactions. We focus on a highly pure sample of $\nu_{\mu}$-CC events between 80 and 560 GeV that interact inside the DeepCore volume. This is the first time that such details of the differential neutrino cross section have been explored at energies between 245 GeV and 1 TeV.
The cross section as a function of inelasticity was parametrized as before [7], and we measured two parameters that define its shape $\langle y\rangle$ and $\lambda$. We then compared the mean inelasticity $\langle y\rangle$ obtained with the expectation from three cross section models and two atmospheric neutrino fluxes.
@@ Acknowledgements.
We thank Yuri Onishchuk for contributions to the early phase of this work. The authors gratefully acknowledge the support from the following agencies and institutions: in the USA, the U.S. National Science Foundation-Office of Polar Programs, U.S. National Science Foundation-Physics Division, U.S. National Science Foundation-EPSCoR, U.S. National Science Foundation-Office of Advanced Cyberinfrastructure, Wisconsin Alumni Research Foundation, Center for High Throughput Computing (CHTC) at the University of Wisconsin–Madison, Open Science Grid (OSG), Partnership to Advance Throughput Computing (PATh), Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS), Frontera and Ranch computing project at the Texas Advanced Computing Center, U.S. Department of Energy-National Energy Research Scientific Computing Center, Particle astrophysics research computing center at the University of Maryland, Institute for Cyber-Enabled Research at Michigan State University, Astroparticle physics computational facility at Marquette University, NVIDIA Corporation, and Google Cloud Platform; in Belgium, Funds for Scientific Research (FRS-FNRS and FWO), FWO Odysseus and Big Science programmes, and Belgian Federal Science Policy Office (Belspo); in Germany, Bundesministerium für Bildung und Forschung (BMBF), Deutsche Forschungsgemeinschaft (DFG), Helmholtz Alliance for Astroparticle Physics (HAP), Initiative and Networking Fund of the Helmholtz Association, Deutsches Elektronen Synchrotron (DESY), and High Performance Computing cluster of the RWTH Aachen; in Sweden, Swedish Research Council, Swedish Polar Research Secretariat, Swedish National Infrastructure for Computing (SNIC), and Knut and Alice Wallenberg Foundation; in the European Union, EGI Advanced Computing for research; in Australia, Australian Research Council; in Canada, Natural Sciences and Engineering Research Council of Canada, Calcul Québec, Compute Ontario, Canada Foundation for Innovation, WestGrid, Digital Research Alliance of Canada, and Arthur B. McDonald Canadian Astroparticle Physics Research Institute; in Denmark, Villum Fonden, Carlsberg Foundation, and European Commission; in New Zealand, Marsden Fund; in Japan, Japan Society for Promotion of Science (JSPS) and Institute for Global Prominent Research (IGPR) of Chiba University; in Korea, National Research Foundation of Korea (NRF); in Switzerland, Swiss National Science Foundation (SNSF).
@@ Appendix A Limitations of the parametrization
The physical interpretation of Eq. (3) requires $y=[0,1]$, which means that $0<$$\langle y\rangle$ $<1$, and $\frac{dp}{dy}$ must be positive, which in turn requires that $\lambda>0$ and $\epsilon>-1$. The allowed physical space and these limits are shown in Fig. 14.
A special case arises when $\lambda<1$, because in that case $\lim_{y\to 0}\frac{dp}{dy}=\infty$. Because of the finite number of events in our data, and also because of presence of kinematic constraints on events with very low inelasticity, the fit might not be able to detect a sharp rise in the distribution near $y=0$. Since there are no data in this region, the integral of the distribution diverges from $1$, thus violating the probability distribution function normalization requirement. This has two undesirable consequences in our analysis. First, this means that the $\langle y\rangle$ variable in Eq. (3) loses its physical interpretation as the mean inelasticity. The second outcome is that breaking of the normalization requirement can cause the minimizer to use $\frac{dp}{dy}$ parameters to modify normalization of the data, which will absorb part of the systematic uncertainties in the study, affecting the correctness of the measurement.
To avert this problem, we adjusted the simulation to data in $\log_{10}\lambda$ instead of $\lambda$ and put a weak half-Gaussian prior on negative $\log_{10}\lambda$ with $\sigma(\log_{10}\lambda)=0.4$ on the negative side. Then, if the data prefer $\lambda<1$, the fit can still allow it with minimal bias, but if this is not the case, the prior will prevent the minimizer from getting stuck in local minima, which is especially relevant when performing a scan to determine confidence regions. The correct behavior of the fit in these cases was confirmed with Monte Carlo tests.